OpenStudy (anonymous):

Is this equation linear, and what is the function for the given x? f(x)=1/2x-11; f(4)

OpenStudy (anonymous):

the equation is not linear. y = 1/2x-11 => (2x-11)y = 1. This is the equation of parabola. f(4) is direct substitution. => f(4) = -1/3